# Trials Factor for Semi-Supervised NN Classifiers in Searches for Narrow Resonances at the LHC

### v2

@@ 1 Introduction
In general, model-independent searches for narrow resonances are of great interest. In recent years the field of particle physics has experienced a growing plethora of anomalous experimental results (for recent reviews see Refs. [29, 30]). Many of them are statistically significant, continue to grow, and remain unexplained by state-of-the-art SM calculations. While some anomalies might eventually find resolution within the SM, persisting ones could contain signs of new physics and may serve as a guide to model building and experimental searches. Of particular relevance to this paper are the multi-lepton anomalies at the LHC [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. These comprise excesses in final states with charged leptons (electrons and muons), missing energy (with and without hadronic and $b$-jets), in processes that have SM Higgs-like topologies. They can be explained by introducing (at least) three additional scalar fields, $H$ and $S$, and $S^{\prime}$ [37, 46]. The scalar $H$ has a mass in the ballpark of 270 GeV, while $S(S^{\prime})$ have masses close to that of the SM Higgs. From the di-lepton component of the multi-lepton anomalies, and assuming $S\rightarrow W^{+}W^{-}\rightarrow\ell\ell\nu\nu,\ell=e,\mu$, the mass of $S$ was determined to be $m_{S}=150\pm 5$ GeV [38], which was later confirmed in Refs. [40, 45]. The first indication of a narrow resonance within this mass range of the di-photon spectrum was reported at $m_{S}=151.5$ GeV in Ref. [47] and further corroborated with more data in Ref. [48]. [^2] At the time of the manuscript submission new excesses at the same mass have been reported in Ref. [49]. Furthermore, an excess around $95$ GeV, denoted $S^{\prime}$ (which acts as the main source of $b$ quarks), has been observed [50, 51, 52, 53, 44, 48]. This provides a compelling reason to seek Higgs-like resonances near the electroweak (EW) scale, an area where SM backgrounds are quite significant. The use of topological constraints through semi-supervision could be crucial in this context, as they help to dampen the impact of these backgrounds, while reducing model dependencies in the search. The $Z\gamma$ final state dataset is therefore selected as an excellent illustrative example for the proposed methodology.
@@ 2.1 Monte Carlo Simulation
The dominant background for our $S(150\,\mathrm{GeV})\rightarrow Z\gamma$ search is the production of a prompt photon in association with a $Z$ boson [54], comprising over $92\%$ of the total background yield [55]. Our simulation accounts for the fact that the $\ell^{+}\ell^{-}\gamma$ invariant mass of 130 GeV to 170 GeV necessitates the $Z$ boson to be off-shell. We configured MadGraph5 aMC@NLO 2.6.7 to include off-shell $Z$ boson contributions, ensuring a comprehensive representation of the physics involved, although virtual-photon contributions were not considered in our simulation.
The event generation utilized next-to-leading-order (NLO) parton distribution functions combined with NLO matrix elements to ensure consistency and accuracy [56]. Generator-level cuts were applied during the event generation in MadGraph5 aMC@NLO 2.6.7, specifically imposing an invariant mass cut on the $\ell\ell\gamma$ system to select events within the 130-170 GeV range. These generated events were then processed through PYTHIA 8.2 [57] for parton showering and Delphes 3 [58] for detector simulation, where reconstructed-level cuts were applied to mimic experimental conditions. Reconstructed-level cuts included detailed pre-selection criteria such as the number of leptons and photons, their transverse momenta, and isolation requirements. Specifically, the $\ell^{+}\ell^{-}\gamma$ candidates were selected by requiring exactly two oppositely charged same-flavor leptons with an invariant mass greater than 40 GeV and one isolated photon with $E_{T}^{\gamma}>25$ GeV.
@@ 5 Conclusion
While the methodology demonstrates that this effect is very small at lower significance values, it is enhanced with decreasing $p$-values (see Figure 7). Despite these complexities, the difference between the baseline with $0\%$ BR cut and $p$-values for higher BR cuts remain limited, showing the method’s proficiency in classifying events. Critically, the results corroborate that semi-supervised NN classifiers introduce controlled fake signals, cementing their viability in HEP analyses. Therefore, this article heralds the precision and utility of semi-supervised NN classifiers for advanced HEP explorations, hinting at their pivotal role in forthcoming discoveries.

### v3

@@ 1 Introduction
In general, model-independent searches for narrow resonances are of great interest. In recent years, the field of particle physics has experienced a growing plethora of anomalous experimental results (for recent reviews, see Refs. [29, 30]). Many of them are statistically significant, continue to grow, and remain unexplained by state-of-the-art SM calculations. While some anomalies might eventually find resolution within the SM, persisting ones could contain signs of new physics and may serve as a guide to model building and experimental searches. Of particular relevance to this paper are the multi-lepton anomalies at the LHC [31, 32, 33, 34]. These can be explained by introducing (at least) two additional scalar fields, $H$ and $S$ [31, 35], where, e.g. $H\rightarrow SS^{(*)}$. The scalar $H$ has a mass in the ballpark of 270 GeV, while the mass of $S$ was determined to be $m_{S}=150\pm 5$ GeV [32]. The first indication of a narrow resonance within this mass range of the di-photon spectrum was reported at $m_{S}=151.5$ GeV in Ref. [36] and further corroborated with more data in Ref. [37].
This provides a compelling reason to seek Higgs-like resonances near the electroweak (EW) scale, an area where SM backgrounds are quite significant. The use of topological constraints through semi-supervision could be crucial in this context, as they help to dampen the impact of these backgrounds, while reducing model dependencies in the search. The $Z\gamma$ final state dataset is therefore selected as an excellent illustrative example for the proposed methodology.
@@ 2.1 Monte Carlo Simulation
The dominant background for our $S(150\,\mathrm{GeV})\rightarrow Z\gamma$ search is the production of a prompt photon in association with a $Z$ boson [38], comprising over $92\%$ of the total background yield [39]. Our simulation accounts for the fact that the $\ell^{+}\ell^{-}\gamma$ invariant mass of 130 GeV to 170 GeV necessitates the $Z$ boson to be off-shell. We configured MadGraph5 aMC@NLO 2.6.7 to include the off-shell $Z$ boson and virtual photon mediated diagram ensuring a comprehensive representation of the physics involved.
The event generation utilized next-to-leading-order (NLO) parton distribution functions, NNPDF 3.0, combined with NLO matrix elements to ensure consistency and accuracy [40]. Generator-level cuts were applied during the event generation in MadGraph5 aMC@NLO 2.6.7, specifically imposing an invariant mass cut on the $\ell\ell\gamma$ system to select events within the 130-170 GeV range. These generated events were then processed through PYTHIA 8.2 [41] for parton showering and Delphes 3 [42] for detector simulation, where reconstructed-level cuts were applied to mimic experimental conditions. Reconstructed-level cuts included detailed pre-selection criteria such as the number of leptons and photons, their transverse momenta, and isolation requirements. Specifically, the $\ell^{+}\ell^{-}\gamma$ candidates were selected by requiring exactly two oppositely charged same-flavor leptons with an invariant mass greater than 40 GeV and one isolated photon with $E_{T}^{\gamma}>25$ GeV.
@@ 5 Conclusion
While the methodology demonstrates that this effect is very small at lower significance values, it is enhanced with decreasing $p$-values (see Figure 7). Despite these complexities, the difference between the baseline with $0\%$ BR cut and $p$-values for higher BR cuts remain limited, showing the method’s proficiency in classifying events. Critically, the results corroborate that semi-supervised NN classifiers introduce controlled fake signals, cementing their viability in HEP analyses. Therefore, this article heralds the precision and utility of semi-supervised NN classifiers for advanced HEP explorations, hinting at their pivotal role in forthcoming discoveries. This work forms a baseline analysis using $Z\gamma$ background processes. Future studies that aim to further this research should include more traditional physics examples and the inclusion of signal processes.